(1 + π‘₯^2) (xdy+ydx) =-2yx^2

1
Expert's answer
2022-02-14T11:38:50-0500

Solution:

The given DE is not solvable, we put dx on right side.

"(1 + 𝑥^2) (xdy+ydx) =-2yx^2 dx
\\ \Rightarrow (1+x^2)(d(xy))=-2yx^2dx"

"\Rightarrow d(xy)=-\dfrac{2yx^2}{1+x^2} dx
\\ \Rightarrow \dfrac{d(xy)}{xy}=-\dfrac{2x}{1+x^2}dx"

On integrating both sides,

"\ln (xy)=-\ln(1+x^2)+\ln C
\\ \Rightarrow \ln (xy(1+x^2))=\ln C
\\ \Rightarrow xy(1+x^2)= C"